OpenStudy (anonymous):

What is the exact volume of a sphere with a diameter of 30 in.? ____ in3 i think the answer is 14137

OpenStudy (anonymous):

@danish071996 @freckles @bohotness

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

u know the formula of volume of sphere ?

OpenStudy (anonymous):

V=4 3πr3

OpenStudy (anonymous):

v=4/3πr3

OpenStudy (anonymous):

@danish071996 ya there?

OpenStudy (anonymous):

its Zayn Malik, and dont remind me D:

OpenStudy (anonymous):

I HAVE NO IDEA WHAT'S THIS

OpenStudy (anonymous):

I cried for 3 hours straight

OpenStudy (anonymous):

I HATE 1D SOOOO IDC

OpenStudy (anonymous):

No comment.

OpenStudy (anonymous):

anyway can you help me @Freez

OpenStudy (anonymous):

SORRY i d k

OpenStudy (campbell_st):

if you know the diameter, find the radius... can you do that

OpenStudy (anonymous):

60 @campbell_st

OpenStudy (campbell_st):

when they ask for an exact value, you need to leave the answer in terms of pi... so the formula becomes \[V = \frac{4 \times r^3}{3} \pi\] just don't multiply by pi

OpenStudy (campbell_st):

you are given the diameter, halve it to get the radius.... so what is half of 30..?

OpenStudy (anonymous):

15

OpenStudy (campbell_st):

great so now calculate \[V = \frac{4 \times 15^3}{3}\] what do you get..?

OpenStudy (anonymous):

4500

OpenStudy (campbell_st):

great so the answer is \[V = 4500 \pi \]

OpenStudy (campbell_st):

just add the units

OpenStudy (anonymous):

thankyou so much